Identify the conic section for the equation. How can you tell the difference between the four conic sections? parabola, ellispe, hyperbola and circle?
Equations:
1. x^2 + y^2 = 36
2. 4x^2-9y^2=36

Answer by fractalier(2101)

1. x^2 + y^2 = 36
This is a circle, centered at (0, 0) with radius 6.
2. 4x^2 - 9y^2 = 36
This is a hyperbola...its proper form is
x^2 / 9 - y^2 / 4 = 1
-----------
Only the parabola has only one variable squared, that's how you identify it.
The hyperbola has the quadratic terms subtracted, that's how you identify it.
The ellipse has the quadratic terms added, with the coefficients of them unequal.
The circle has the quadratic terms added, with the coefficients of them equal.